Please answer 3 - 5 thank you!3. In his neighborhood, a businessman would be setting up a repair shop. At a cash price of P 2.8 Million, he would be gettin - 5 thank you!
Answer #1

EUAC for Plant 1 = 2 billion + 2 milion(A/P, 8%, 15) + 3 million (A/P, 8%, 15)

(A/P, 8%, 15) = 0.1168 (using factor table for 8% and 15 years)

EUAC = 2 billion + 2 million *0.1168 + 3 * 0.1168 = 2 billion + 233600 + 350400 = 2,000,584,000

Pcost of plant after 15 years = 1.5 billion(P/A, *%, 15) = 12839250000

EUAC for plant 2 = 1.25 billion + 12839250000(A/P, 8%, 15) + 1.5 million (A/P, 8%, 15) + 3 million (A/P, 8%, 15)

(P/A, 8%, 15) = 8.5595 (from factor table)

EUAC = 1.25 billion + 1,499,624,400 + 175200 + 350400 = 3,000,150,000

SInce, Option 1 has lower EUAC, It is better option than option 2.